OpenStudy (anonymous):

y=sin^-1(2x/1+x^2) differentite

OpenStudy (rogue):

The derivative of the arcsine function is as follows\[\frac {d}{dx} \left[ \sin^{-1} u \right] = \frac {u'}{\sqrt {1 - u^2}}\] Use that to find the derivative :)

hartnn (hartnn):

ther's a shortcut, put x= tan t

OpenStudy (anonymous):

ok i try

OpenStudy (anonymous):

cant we do it directly by using inverse trinometri derivative

hartnn (hartnn):

algebraically cumbersome,putting x=tan t will simplify this greatly.

OpenStudy (rogue):

Hmm, Hartnn, how would we do it with that substitution? Would it require implicit differentiation?

OpenStudy (anonymous):

its sin^-1tan2tan^-1x what to do next

hartnn (hartnn):

see, x=tan t means 2x/(1+x^2)=sin 2t so sin^-1 sin 2t=2t=y so y is just 2 tan^-1 x so that u can use standard formula for tan^-1 x got it?

OpenStudy (anonymous):

2x/(1+x^2)=sin 2t from whre it came

hartnn (hartnn):

2tan t /(1+tan^2 t)=2 tan t/sec^2 t=2sin t cos^2t/cos t=2 sin t cos t=sin 2t .........got it?